Rationalize the denominator and simplify

6
________________
(3 √7 – 2 √6) (3 √7 + 2 √6)

Answer by edjones(8007)

6
________________
(3 √7 – 2 √6) (3 √7 + 2 √6)
.
Let a=3sqrt(7) and b=2sqrt(6)
=6/((a-b)(a+b))
=6/(a^2-b^2)
=6/(63-24)
=6/39
=2/13
.
Ed